OpenStudy (anonymous):

Joe wrote a book report. The report was 10 pages long. Jake's book report was 1/5 as long as Joe's report. How long was Jake's report? Can you please explain how to reach the answer?

OpenStudy (anonymous):

If Jake's report is 1/5 as long as Joe's it is 1/5 times the total length of Joe's report. 1/5*10=10/5=2. since it is 1/5 as long if you put 5 of them together you would get Joe's report. 2*5=10. So the solution is 2 pages.